OpenStudy (anonymous):

Find the side of a square whose diagonal is of the given measure. Given = 12 sqrt 10ft

OpenStudy (anonymous):



OpenStudy (anonymous):

so \[a \sqrt{2}=12\sqrt{10}\]

OpenStudy (anonymous):

\[a=12\sqrt{10}\div \sqrt{2}\]

OpenStudy (anonymous):

a=\[12\sqrt{5}\]